OpenStudy (anonymous):

Solve |x + 6| = 12. 6 -18 {6, -18} no solution

OpenStudy (anonymous):

try put the values of answers

OpenStudy (mertsj):

|x+6|=12 means that x+6 is exactly 12 units from 0. That means it could be 12 or -12 since both are 12 units from 0. So we write: x+6=12 or x+6=-12 Now solve each part.